# What Expressivity Theory Misses: Message Passing Complexity for GNNs

### v1

@@ C.6 Complexity bounds in §˜5
$\text{For CIN \& FragNet:}\>\>O(\log(sr))\quad\quad\text{For GSN:}\>\>\leq s/2\log(r+1)\quad\quad\text{For all others:}\>\>\geq s\log(r)$
@@
FragNet: Notice that for FragNet $L=2$ is the minimal depth required to solve this task. With $L=2$ a sufficient set of messages involves a message from $u$ to the fragment node that represents this cycle (with message success probability $1/s$) and from the fragment node to $u$ (with message success probability $1/(r+2)$). Using Lemma˜C.10, the bound follows immediately.
CIN: Again, for CIN $L=2$ is the minimal depth required to solve this task. A sufficient set of messages involves a message from the ring node (which contains as initial encoding all node features from the nodes in the ring) to an edge node. And from the edge node to $v$. Notice that the edge node has degree $s+2$ (an edge to every other edge node in the cycle, an edge to the ring node and two edges to the original nodes) and $v$ has degree $2r+1$ (edges to all $r$ neighboring nodes in G{G}, edges to corresponding edge nodes, and self-loop). Hence, we have message success probabilities $1/(s+2)$ and $1/(2r+1)$ Again, using Lemma˜C.10, the bound follows immediately.
GSN: Notice that for GSN the cycle information is already contained in the initial node-feature. So the task is simply to propagate the node feature information to $v$ for a distance of at most $s/2$. Hence, a sufficient set of messages consists of the messages on a path of length $L\leq s/2$ from $u$ to $v$, each with success probability $1/(r+1)$. Again, using Lemma˜C.10, the bound follows immediately.
@@
Let $\mathcal{G}$} be a family of graphs with unique node features, i.e., for any ${G}\in{\mathcal{G}}$, no two nodes $v,w\in{\mathcal{V}_{G}}$ have the same node features. Then the WL output after one iteration $\{\mskip-5.0mu\{{\text{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}\}\mskip-5.0mu\}$ differs for any two non-isomorphic graphs in $\mathcal{G}$}.
@@
Let ${G}_{1},{G}_{2}\in{\mathcal{G}}$ be two non-isomorphic graphs. Suppose, for contradiction, that the WL output after one iteration is the same for both graphs, i.e., $\{\mskip-5.0mu\{{\text{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}_{1}\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\text{WL}}_{w}^{1}\mid w\in{\mathcal{V}_{G}}_{2}\}\mskip-5.0mu\}$ as multisets.
Since all node features in ${G}_{1}$ and ${G}_{2}$ are unique within each graph, the initial WL labels ${\text{WL}}_{v}^{0}$ are distinct for all nodes $v$ in each graph. After one iteration, each ${\text{WL}}_{v}^{1}$ consists of the node’s original feature combined with the multiset of its neighbors’ features, i.e., ${\text{WL}}_{v}^{1}=\left({\text{WL}}_{v}^{0},\{\mskip-5.0mu\{{\text{WL}}_{u}^{0}\mid u\in{N}(v)\}\mskip-5.0mu\}\right)$.
Given that $\{\mskip-5.0mu\{{\text{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}_{1}\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\text{WL}}_{w}^{1}\mid w\in{\mathcal{V}_{G}}_{2}\}\mskip-5.0mu\}$, there must exist a bijection $f:{\mathcal{V}_{G}}_{1}\rightarrow{\mathcal{V}_{G}}_{2}$ such that ${\text{WL}}_{v}^{1}={\text{WL}}_{f(v)}^{1}$ for all $v\in{\mathcal{V}_{G}}_{1}$.
This implies that for each $v\in{\mathcal{V}_{G}}_{1}$: (1) ${\text{WL}}_{v}^{0}={\text{WL}}_{f(v)}^{0}$ (the original node features match), and (2) $\{\mskip-5.0mu\{{\text{WL}}_{u}^{0}\mid u\in{N}(v)\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\text{WL}}_{w}^{0}\mid w\in{N}(f(v))\}\mskip-5.0mu\}$ (the multisets of neighbor features match).
@@
From (2), for each $v\in{\mathcal{V}_{G}}_{1}$, there must exist a bijection $g_{v}:{N}(v)\rightarrow{N}(f(v))$ such that ${\text{WL}}_{u}^{0}={\text{WL}}_{g_{v}(u)}^{0}$ for all $u\in{N}(v)$.
@@ Appendix D Additional results and Experimental Details
  • Validation on Erdős-Rényi (ER) graphs, confirming that MPC accurately matches performance trends also on a different graph family.
  • Real-world validation on graphs from ZINC Dwivedi et al. [2023] and peptides datasets Dwivedi et al. [2022] showing that MPC explains performance better than classical expressivity theory for real-world graph structures.
@@ D.1 Implementation Details and Experimental Setup
We use pytorch Paszke et al. [2019] and torch-geometric Fey and Lenssen [2019] (released under an MIT license) for the implementation of all models. For optimization, we use the Adam optimizer Kingma and Ba [2015]. The training, validation and test sets containing random $r$-regular graphs are generated using networkx Hagberg et al. [2008]. For all standard MPNNs we use their pytorch geometric implementation. Note that GSN Bouritsas et al. [2023] only specifies the additional node features and not the downstream MPNN. We use the most common MPNN, GCN, as downstream MPNN for all experiments for GSN. For CIN Bodnar et al. [2021a] we use our own implementation following exactly the method proposed in their paper. Additionally, for FragNet, we use a custom implementation of their FR-WL model without edge representations. For CIN, GSN, and FragNet we use a fragmentation scheme identifying every cycle of size at most 6 (unless otherwise noted). For all models we use an initial feature embedding layer and a final output MLP. Additionally, we use BatchNorm Ioffe and Szegedy [2015] for normalization. We found little difference between learning rates in $\{0.001,0.005,0.01\}$ for all models and tasks. The shown results are for the learning rate $0.005$. We train all models for all settings using three different seeds for a maximum of 50 epochs, showing the average results. All experiments are conducted on NVIDIA GeForce GTX 1080 GPUs with 16GB memory allocation per job. Training times vary by model architecture, ranging from 10-30 minutes. Monte Carlo simulations for complexity calculations run on a single Intel Xeon E5-2630 v4 CPU (2.20GHz) and complete in under 10 seconds. The peptides dataset of the lrgb Dwivedi et al. [2022] is released under a CC BY-NC 4.0 license. The ZINC dataset Dwivedi et al. [2023] is distributed under a custom license (free to use for everyone).
@@ D.3 Propagating Information
Figure 17: MPC captures task difficulty on real-world graphs. Test accuracy by distance for the information propagation task $f_{v}({G})={\bm{X}}_{u}$ across varying training dataset sizes, with simulated MPC complexity shown for real-world graphs from the peptides-func/peptides-struct dataset Dwivedi et al. [2022]. Higher MPC complexity corresponds to greater learning difficulty, as evidenced by increased sample complexity requirements, demonstrating that MPC effectively captures practical task difficulty beyond synthetic graph families.
Additionally, we repeat the information propagation task on graphs from the real-world peptides-func/peptides-struct dataset. It comprises larger peptide molecules that require models to consider long-range interactions Dwivedi et al. [2022]. Therefore, the information propagation task, testing the models’ ability to exchange information over varying distances, is particularly well-suited for this dataset.
@@ D.4 Extracting topological information
We additionally test the ring transfer capabilities of the models on real-world graphs from the ZINC molecular regression dataset Dwivedi et al. [2023]. The ZINC dataset comprises small molecules and the benchmark task is to predict the penalized logP score which involves the number of cycles. Therefore the graphs from the ZINC dataset are well-suited to test our models cycle detection capabilities. For this, we first filter all graphs to contain rings of size $s$. Second, we randomly choose a node $v$ that is part of cycle $s$ for each graph. The labels are assigned in the same way as for the random regular graphs and the ER graphs.

### v2

@@ C.6 Complexity bounds in §˜5
$\text{For CIN \& FragNet:}\>\>O(\log(sr))\quad\quad\text{For GSN:}\>\>\leq\lceil s/2\rceil\log(r+1)\quad\quad\text{For all others:}\>\>\geq s\log(r)$
@@
FragNet: Notice that for FragNet $L=2$ is the minimal depth required to solve this task. With $L=2$ a sufficient set of messages involves a message from $u$ to the fragment node that represents this cycle (with message success probability $1/s$) and from the fragment node to $v$ (with message success probability $1/(r+2)$). Using Lemma˜C.10, the bound follows immediately.
CIN: Again, for CIN $L=2$ is the minimal depth required to solve this task. A sufficient set of messages involves a message from the ring node (which contains as initial encoding all node features from the nodes in the ring) to an edge node. And from the edge node to $v$. Notice that the edge node has degree $s+2$ (an edge to every other edge node in the cycle, an edge to the ring node and two edges to the original nodes) and $v$ has degree $2r+1$ (edges to all $r$ neighboring nodes in G{G}, edges to corresponding edge nodes, and self-loop). Hence, we have message success probabilities $1/(s+2)$ and $1/(2r+1)$. Again, using Lemma˜C.10, the bound follows immediately.
GSN: Notice that for GSN the cycle information is already contained in the initial node-feature. So the task is simply to propagate the node feature information to $v$ for a distance of at most $\lceil s/2\rceil$. Hence, a sufficient set of messages consists of the messages on a path of length $L\leq\lceil s/2\rceil$ from $u$ to $v$, each with success probability $1/(r+1)$. Again, using Lemma˜C.10, the bound follows immediately.
@@
Let $\mathcal{G}$} be a family of graphs with unique node features, i.e., for any ${G}\in{\mathcal{G}}$, no two nodes $v,w\in{\mathcal{V}_{G}}$ have the same node features. Then the WL output after one iteration $\{\mskip-5.0mu\{{\textup{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}\}\mskip-5.0mu\}$ differs for any two non-isomorphic graphs in $\mathcal{G}$}.
@@
Let ${G}_{1},{G}_{2}\in{\mathcal{G}}$ be two non-isomorphic graphs. Suppose, for contradiction, that the WL output after one iteration is the same for both graphs, i.e., $\{\mskip-5.0mu\{{\textup{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}_{1}\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\textup{WL}}_{w}^{1}\mid w\in{\mathcal{V}_{G}}_{2}\}\mskip-5.0mu\}$ as multisets.
Since all node features in ${G}_{1}$ and ${G}_{2}$ are unique within each graph, the initial WL labels ${\textup{WL}}_{v}^{0}$ are distinct for all nodes $v$ in each graph. After one iteration, each ${\textup{WL}}_{v}^{1}$ consists of the node’s original feature combined with the multiset of its neighbors’ features, i.e., ${\textup{WL}}_{v}^{1}=\left({\textup{WL}}_{v}^{0},\{\mskip-5.0mu\{{\textup{WL}}_{u}^{0}\mid u\in{N}(v)\}\mskip-5.0mu\}\right)$.
Given that $\{\mskip-5.0mu\{{\textup{WL}}_{v}^{1}\mid v\in{\mathcal{V}_{G}}_{1}\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\textup{WL}}_{w}^{1}\mid w\in{\mathcal{V}_{G}}_{2}\}\mskip-5.0mu\}$, there must exist a bijection $f:{\mathcal{V}_{G}}_{1}\rightarrow{\mathcal{V}_{G}}_{2}$ such that ${\textup{WL}}_{v}^{1}={\textup{WL}}_{f(v)}^{1}$ for all $v\in{\mathcal{V}_{G}}_{1}$.
This implies that for each $v\in{\mathcal{V}_{G}}_{1}$: (1) ${\textup{WL}}_{v}^{0}={\textup{WL}}_{f(v)}^{0}$ (the original node features match), and (2) $\{\mskip-5.0mu\{{\textup{WL}}_{u}^{0}\mid u\in{N}(v)\}\mskip-5.0mu\}=\{\mskip-5.0mu\{{\textup{WL}}_{w}^{0}\mid w\in{N}(f(v))\}\mskip-5.0mu\}$ (the multisets of neighbor features match).
@@
From (2), for each $v\in{\mathcal{V}_{G}}_{1}$, there must exist a bijection $g_{v}:{N}(v)\rightarrow{N}(f(v))$ such that ${\textup{WL}}_{u}^{0}={\textup{WL}}_{g_{v}(u)}^{0}$ for all $u\in{N}(v)$.
@@ Appendix D Additional results and Experimental Details
  • Validation on Erdős-Rényi (ER) graphs, confirming that MPC accurately matches performance trends also on a different synthetic graph family.
  • Real-world validation on graphs from ZINC [15] and peptides datasets [14] showing that MPC explains performance better than classical expressivity theory for real-world graph structures.
  • Results showing that insight on synthetic proxy tasks can mirror performance trends on real-world benchmarks where the precise tasks are unknown.
@@ D.1 Implementation Details and Experimental Setup
We use pytorch [39] and torch-geometric [17] (released under an MIT license) for the implementation of all models. For optimization, we use the Adam optimizer [30]. The training, validation and test sets containing random $r$-regular graphs are generated using networkx [24]. For all standard MPNNs we use their pytorch geometric implementation. Note that GSN [8] only specifies the additional node features and not the downstream MPNN. We use the most common MPNN, GCN, as downstream MPNN for all experiments for GSN. For CIN [5] we use our own implementation following exactly the method proposed in their paper. Additionally, for FragNet, we use a custom implementation of their FR-WL model without edge representations using a GCN as base MPNN. For CIN, GSN, and FragNet we use a fragmentation scheme identifying every cycle of size at most 6 (unless otherwise noted). For all models we use an initial feature embedding layer and a final output MLP. Additionally, we use BatchNorm [26] for normalization. We found little difference between learning rates in $\{0.001,0.005,0.01\}$ for all models and tasks. The shown results are for the learning rate $0.005$. We train all models for all settings using three different seeds for a maximum of 50 epochs, showing the average results. All experiments are conducted on NVIDIA GeForce GTX 1080 GPUs with 16GB memory allocation per job. Training times vary by model architecture, ranging from 10-30 minutes. Monte Carlo simulations for complexity calculations run on a single Intel Xeon E5-2630 v4 CPU (2.20GHz) and complete in under 10 seconds. The peptides dataset of the lrgb [14] is released under a CC BY-NC 4.0 license. The ZINC dataset [15] is distributed under a custom license (free to use for everyone).
@@ D.3 Propagating Information
Figure 17: MPC captures task difficulty on real-world graphs. Test accuracy by distance for the information propagation task $f_{v}({G})={\bm{X}}_{u}$ across varying training dataset sizes, with simulated MPC complexity shown for real-world graphs from the peptides-func/peptides-struct dataset [14]. Higher MPC complexity corresponds to greater learning difficulty, as evidenced by increased sample complexity requirements, demonstrating that MPC effectively captures practical task difficulty beyond synthetic graph families.
Additionally, we repeat the information propagation task on graphs from the real-world peptides-func/peptides-struct dataset. It comprises larger peptide molecules that require models to consider long-range interactions [14]. Therefore, the information propagation task, testing the models’ ability to exchange information over varying distances, is particularly well-suited for this dataset.
@@ D.4 Extracting topological information
We additionally test the ring transfer capabilities of the models on real-world graphs from the ZINC molecular regression dataset [15]. The ZINC dataset comprises small molecules and the benchmark task is to predict the penalized logP score which involves the number of cycles. Therefore the graphs from the ZINC dataset are well-suited to test our models cycle detection capabilities. For this, we first filter all graphs to contain rings of size $s$. Second, we randomly choose a node $v$ that is part of cycle $s$ for each graph. The labels are assigned in the same way as for the random regular graphs and the ER graphs.